# 70% Size, 100% Accuracy: Lossless LLM Compression for Efficient GPU Inference via Dynamic-Length Float

### v1

@@ 2 Motivation: Is Lossless Compression of LLMs Worth Studying?
Much of the motivation behind our work lies in understanding whether lossless compression of LLMs, which preserves 100% identical output behavior compared to the original uncompressed model, is a practical direction worthy of further study. Specifically, how does DFloat11, which compresses LLMs to approximately 11 bits, compare to widely used lossy quantization techniques [7, 21], where models are typically reduced to even lower bit-widths (e.g., 8-bit or 4-bit)?
@@
That being said, the argument that “current benchmarks fail to capture the performance gap between 8-bit compressed and 16-bit uncompressed models” is itself constrained by the limitations of the current benchmarking landscape, making it difficult to produce abundant supporting evidence. Nonetheless, some reports have begun to highlight such gaps. For example, human evaluations on LLM Arena [^1] https://x.com/lmarena_ai/status/1835760196758728898 show a notable performance drop between Llama-3.1-405B-Instruct [11] and its 8-bit counterpart (Llama-3.1-405B-Instruct-FP8), particularly under coding (1293 vs. 1277) and long-query (1282 vs. 1275) tasks. Similarly, quantizing DeepSeek-R1-Distill-Llama-70B [12] from 16 bits to 8 bits results in a 23.7% drop on GPQA (from 9.51% to 7.25%). [^2] https://huggingface.co/RedHatAI/DeepSeek-R1-Distill-Llama-70B-quantized.w8a8 Furthermore, reasoning, a core capability of modern LLMs, appears especially sensitive to compression loss. Recent benchmark [23] reveals that quantizing DeepSeek-R1-Distill-Qwen-1.5B with 8-bit SmoothQuant [34] (for weight, attention, and KV cache) leads to an average 9.09% drop in reasoning tasks (48.82% to 44.29%) across datasets like AIME, MATH-500, GPQA-Diamond, and LiveCodeBench. We leave more evidence exploring the performance gap between 8-bit quantized and uncompressed model in Appendix D.
@@
To eliminate this burden, DFloat11 offers a compelling alternative: delivering 100% identical performance to the original model, while consuming only $\sim$70of the memory footprint with many throughput benefits, which is a unique and practical offering for resource-constrained deployment settings.
3 Method
In this section, we present our framework for lossless compression of LLM weights. We begin with background on BFloat16, entropy coding, and the GPU computation and memory model. Next, we motivate our approach by analyzing the information efficiency of BFloat16 in representing model weights. We then introduce our proposed floating-point format, Dynamic-Length Float (DFloat11), along with its custom decompression kernel designed for efficient GPU inference.
3.1 Preliminary
Brain Float (BFloat16)
The weights of large language models (LLMs) are typically represented using floating-point number formats. Recent state-of-the-art LLMs predominantly employ the 16-bit Brain Floating Point format (BFloat16 or BF16), which balances numerical precision and memory efficiency. BF16 allocates its 16 bits as follows: 1 sign bit, 8 exponent bits, and 7 mantissa bits. The numerical value represented by a BF16 number is computed as:
where $\mathrm{mantissa}$ is interpreted as a binary fractional value. Compared to FP32, BF16 uses half the bytes to represent the same number of parameters while having a similar numerical range, due to having the same number of exponent bits. In contrast to FP16, BF16 offers a wider numerical range, reducing the risk of overflow during model training and inference.
Entropy Coding
Entropy coding is a core technique in lossless data compression that leverages statistical redundancy to reduce data size. Several widely used methods fall under this category, including Huffman coding [17], arithmetic coding [20], and Asymmetric Numeral Systems (ANS) [5]. Among these, Huffman coding is one of the earliest and most widely adopted, which uses variable-length encoding to minimize the size of encoded data. It assigns shorter binary codes to more frequent symbols and longer codes to less frequent ones. The codes are decoded using a prefix-free binary tree, known as a Huffman tree. Due to the prefix-free property of Huffman codes, no code is a prefix of any other, which ensures unique decodability of the encoded bitstream without the need for delimiters. The tree is constructed based on symbol frequencies and is provably optimal for any given frequency distribution. Huffman coding is widely used in file compression and data transmission. However, decoding in a massively parallel manner remains challenging due to its inherently sequential nature.
GPU Computation and Memory Paradigm
GPUs are designed to perform computations in a massively parallel manner. A modern GPU consists of thousands of threads, which are organized into blocks and executed on streaming multiprocessors (SMs). Each block of threads has access to a small, fast on-chip memory known as shared memory or SRAM. This memory offers low latency and high throughput, making it much more efficient for frequent read and write operations compared to the global high-bandwidth memory (HBM), which is off-chip and shared across all SMs. The capacity of shared memory is limited, typically having up to 100 KB per block. In this work, we leverage the fast access characteristics of SRAM to enable efficient on-the-fly decompression of losslessly compressed weights during inference.
Figure 1: (Left) The allocation of bits for the components of BFloat16. (Right 3) The information content, as measured by Shannon Entropy, of the components (sign, exponent, mantissa) of BFloat16 weights in various LLMs.
3.2 Technical Motivation: BFloat16 Representation is Information Inefficient
To motivate the lossless compression of LLM weights, we analyze the compressibility of the BFloat16 components (sign, exponent, and mantissa) in the weights of recent large language models. Specifically, we use Shannon entropy to quantify the information content of all parameters within the linear projection matrices of an LLM. The Shannon entropy $H(\cdot)$ is defined as:
$H(X)=-\sum_{x\in\mathcal{X}}p(x)\log_{2}p(x)$ (2)
where $X$ is a discrete random variable with support $\mathcal{X}$, and $p:\mathcal{X}\to[0,1]$ denotes its probability mass function. We present the computed entropy values in Figure 1. As shown, the entropy of the sign and mantissa components is close to their respective bit widths, indicating limited potential for compression. In contrast, the exponent exhibits significantly lower entropy, approximately 2.6 bits versus its allocated 8 bits, suggesting substantial opportunities for lossless compression.
To better understand this discrepancy, we visualize the relative frequency for all values of BFloat16 components in Figure 7 in the Appendix, and plot the ranked frequency distribution of exponent values in Figure 8 in the Appendix. As shown, the sign and mantissa values are distributed relatively uniformly across the entire value ranges. However, the distribution of exponent values is highly imbalanced: out of the 256 possible 8-bit values, only around 40 are used, with the rest never appearing in the weights. Moreover, the ranked frequencies decay rapidly towards zero. These findings explain the low entropy of the exponent and highlight the potential for significant compression. In the following sections, we leverage the low information content of BFloat16 exponents to design a lossless compression framework that enables efficient GPU inference.
Figure 2: An illustration of our proposed format Dynamic-Length Float for compressing the BFloat16 weights of LLMs losslessly.
3.3 Dynamic-Length Float: A Lossless LLM Compression Framework for Efficient GPU Inference
To address the substantial information inefficiency in the BFloat16 representation of LLM weights, we propose a lossless compression framework that encodes floating-point parameters using entropy coding. Specifically, we build a Huffman tree based on the distribution of exponents from all BFloat16 weights within the linear projection matrices of an LLM. We then compress the exponents using Huffman coding, while preserving the original signs and mantissas. Exponents are encoded and tightly bit-packed into a byte array, $\mathbf{EncodedExponent}$, while the sign and mantissa are left uncompressed and stored in a separate byte array $\mathbf{PackedSignMantissa}$. Figure 2 illustrates Dynamic-Length Float (DFloat11 or DF11), our proposed format for compactly representing floating-point model parameters.
The Core Challenge: Efficient GPU Inference with Compressed Weights
Although Dynamic-Length Float enables effective lossless compression of LLMs, a key challenge remains: how to perform efficient GPU inference with these compressed weights. Entropy-coded weights, due to their variable-length encoding, cannot be directly used in matrix multiplications. Therefore, each weight matrix is decompressed on-the-fly to its original BFloat16 representation when needed during inference, and immediately discarded after the matrix multiplication completes to save memory. Traditional Huffman decoding involves bit-by-bit traversal of the Huffman tree to decode each element, a process that is inherently sequential and fits poorly with the massively parallel execution model of GPUs. Naively assigning a single thread to decode each value results in severe underutilization of GPU resources and high latency. Overcoming this bottleneck is critical for making compressed inference practical.
In the following paragraphs, we present our solution in detail: a set of hardware-aware algorithmic designs tailored for low-latency decoding of entropy-coded weights in a massively parallel manner. Our approach consists of three key components:
  1. 1.
    Decomposing a monolithic prefix-free lookup table (LUT) into multiple compact LUTs that fit within GPU SRAM;
  2. 2.
    Introducing a two-phase kernel design that leverages lightweight auxiliary variables to efficiently coordinate read/write operations of threads;
  3. 3.
    Performing decompression at the transformer block level to boost throughput and minimize latency.
@@ 3.3 Dynamic-Length Float: A Lossless LLM Compression Framework for Efficient GPU Inference
  • $\mathbf{EncodedExponent},\mathbf{PackedSignMantissa}$: byte arrays
  • $\mathbf{LUT}_{1},\mathbf{LUT}_{2},\mathbf{LUT}_{3},\mathbf{LUT}_{4},\mathbf{CodeLengths}$: 8-bit unsigned integer arrays of size 256
  • $\mathbf{Gaps}$: 5-bit unsigned integer array (one entry per thread in each block)
  • $\mathbf{BlockOutputPos}$: 32-bit unsigned integer array (one entry per block)
  • $\mathbf{Outputs}$: BFloat16 array, for storing results
  • $B,T,n,R$: the number of blocks, number of threads, number of bytes in $\mathbf{EncodedExponent}$ processed by each thread, a reserved value, respectively
3:     Divide $\mathbf{EncodedExponent}$ into chunks:
4:         $\mathbf{EncodedExponent}_{1},\dots,\mathbf{EncodedExponent}_{B}$ of size $nT$ bytes each
@@
6:         Load $\mathbf{EncodedExponent}_{b}$ into SRAM
7:         Divide $\mathbf{EncodedExponent}_{b}$ into chunks:
8:             $\mathbf{EncodedExponent}_{b,1},\dots,\mathbf{EncodedExponent}_{b,T}$ of size $n$ bytes each
9:         Load $\mathbf{LUT}_{1},\mathbf{LUT}_{2},\mathbf{LUT}_{3},\mathbf{LUT}_{4},\mathbf{CodeLengths}$ into SRAM
10:         Initialize integer arrays $\mathbf{NumElements}[1\dots T],\mathbf{ThreadOutputPos}[1\dots T]$ with all 00s
11:         for all $t\leftarrow 1,\dots,T$ (in parallel across threads) do
12$\triangleright$ Phase 1: Each thread determines its initial output position
13:              $\mathrm{BitOffset}\leftarrow\mathbf{Gaps}[bT+t]$
14:              while $\mathrm{BitOffset}<8n$ do
15:                  Read the next 4 bytes of $\mathbf{EncodedExponent}_{b,t}$, starting from the $\mathrm{BitOffset}$-th bit, into $\mathrm{Byte}_{1\dots 4}$
16:                  $\mathrm{Exponent}\leftarrow R$
@@
18:                  while $\mathrm{Exponent}=R$ do
19:                       $\mathrm{Exponent}\leftarrow\mathbf{LUT}_{i}[\mathrm{Byte}_{i}]$
20:                       $i\leftarrow i+1$
21:                  end while
22:                  $\mathrm{BitOffset}\leftarrow\mathrm{BitOffset}+\mathbf{CodeLengths}[\mathrm{Exponent}]$
23:                  $\mathbf{NumElements}[t]\leftarrow\mathbf{NumElements}[t]+1$
24:              end while
25:              Thread Synchronization Barrier
26:              $\mathbf{ThreadOutputPos}[t]\leftarrow\mathbf{BlockOutputPos}[b]+\sum_{i=1}^{t-1}\mathbf{NumElements}[i]$
27$\triangleright$ Phase 2: Writing decoded BFloat16s to the appropriate positions
28:              $\mathrm{BitOffset}\leftarrow\mathbf{Gaps}[bT+t]$
29:              while $\mathrm{BitOffset}<8n$ do
30:                  Read the next 4 bytes of $\mathbf{EncodedExponent}_{b,t}$, starting from the $\mathrm{BitOffset}$-th bit, into $\mathrm{Byte}_{1\dots 4}$
31:                  $\mathrm{Exponent}\leftarrow R$
32:                  $i\leftarrow 1$
33:                  while $\mathrm{Exponent}=R$ do
34:                       $\mathrm{Exponent}\leftarrow\mathbf{LUT}_{i}[\mathrm{Byte}_{i}]$
35:                       $i\leftarrow i+1$
36:                  end while
37:                  $\mathrm{Byte}\leftarrow\mathbf{PackedSignMantissa}\big{[}\mathbf{ThreadOutputPos}[t]\big{]}$

### v2

@@ Appendix A Discussion: Is Quantization a Universal Solution?
Much of the motivation behind our work lies in understanding whether lossless compression of large-scale models such as LLMs, which preserves 100% identical output behavior compared to the original uncompressed model, is a practical direction worthy of further study. Specifically, how does DFloat11, which compresses LLMs to approximately 11 bits, compare to widely used lossy quantization techniques [14, 32], where models are typically reduced to even lower bit-widths (e.g., 8-bit or 4-bit)?
@@
That being said, the argument that “current benchmarks fail to capture the performance gap between 8-bit compressed and 16-bit uncompressed models” is itself constrained by the limitations of the current benchmarking landscape, making it difficult to produce abundant supporting evidence. Nonetheless, some reports have begun to highlight such gaps. For example, human evaluations on LLM Arena [^1] https://x.com/lmarena_ai/status/1835760196758728898 show a notable performance drop between Llama-3.1-405B-Instruct [18] and its W8A8 counterpart (Llama-3.1-405B-Instruct-FP8), particularly under coding (1293 vs. 1277) and long-query (1282 vs. 1275) tasks. Similarly, quantizing DeepSeek-R1-Distill-Llama-70B [19] from 16 bits to 8 bits results in a 23.7% drop on GPQA (from 9.51% to 7.25%). [^2] https://huggingface.co/RedHatAI/DeepSeek-R1-Distill-Llama-70B-quantized.w8a8 Furthermore, reasoning, a core capability of modern LLMs, appears especially sensitive to compression loss. Recent benchmark [34] reveals that quantizing DeepSeek-R1-Distill-Qwen-1.5B with 8-bit SmoothQuant [43] (for weight, attention, and KV cache) leads to an average 9.09% drop in reasoning tasks (48.82% to 44.29%) across datasets like AIME, MATH-500, GPQA-Diamond, and LiveCodeBench. We leave more evidence exploring the performance gap between 8-bit quantized and uncompressed model in Appendix H.
@@
To eliminate this burden, DFloat11 offers a compelling alternative: delivering 100% identical performance to the original model, while consuming only $\sim$70% of the memory footprint with many throughput benefits, which is a unique and practical offering for resource-constrained deployment settings.
Appendix B Extended Related Works
Data Formats for Model Weights
Compressed data formats typically aim for lower bit-widths. For example, FP8—which comes in both E4M3 (4 exponent bits, 3 mantissa bits, plus 1 sign bit) and E5M2 configurations—has seen reasonable adoption in LLM training and development. Integer formats like INT8 have also been well explored, as in LLM.int8() [7] and its following works. Formats with a stronger emphasis on efficiency, such as FP4, INT4, NF4 [8], and AF4 [49], use only 4 bits. In this work, we primarily focus on formats with $\geq$8 bits, as benchmark literature [46, 17, 34] often suggests that 8-bit quantization results in negligible performance drop—though we show in Section A that this claim is likely skewed due to evaluation selectiveness and benchmark limitations.
Lossless Model Compression
While lossy model compression techniques such as pruning and quantization [13, 32, 14] have received widespread attention, lossless model compression remains a relatively underexplored area. Upon careful investigation, we identified roughly four prior works that have made meaningful efforts in this space. Deep Compression [20] is a foundational work, applying Huffman coding [24] to quantized CNN models and achieving an additional $\sim$22% compression gain for model checkpoints. ZipNN [23] extended this idea to language models, comparing its results to classic lossless compression tools such as zlib [9] and zstd [^3] https://github.com/facebook/zstd and demonstrated superior compression gains. However, this line of work—including their industry counterparts, such as ezm7444https://github.com/liuliu/s4nnc/pull/11 https://encode.su/threads/4067-Good-Compressors-for-16-bit-floats—is limited in that its efficiency gains only apply to storage (reducing the size of model checkpoints) but offer no benefits during inference. While such storage savings are meaningful in large-scale training settings—where frequent snapshotting and checkpoint rollbacks are needed [41]—they have limited impact for everyday LLM end-users. Model downloading is typically a one-time cost, so even if a model checkpoint is compressed by 50%, it only cuts the download time at most by half, presumably over the model’s entire lifecycle of deployment. Furthermore, checkpoints are usually stored on disk, where terabytes of capacity are easily available, making up a much looser constraint compared to GPU HBM (High Bandwidth Memory); one of the main resource constraints during inference.
Appendix C Frequency Distribution of BFloat16 Values
Figure 8 presents the frequency distribution for distinct values of sign, exponent, and mantissa bits in the BFloat16 weights of LLMs. Figure 9 shows the sorted frequency of exponent values of LLM weights.
Appendix D Pseudo-code of the GPU kernel for DFloat11 Decompression
Algorithm 1 presents the pseudo-code of the two-phase GPU kernel for decompressing DFloat11 to BFloat16.
@@ Appendix D Pseudo-code of the GPU kernel for DFloat11 Decompression
  • $\mathsf{EncodedExponent},\mathsf{PackedSignMantissa}$: byte arrays
  • $\mathsf{LUT}_{1},\dots,\mathsf{LUT}_{k},\mathsf{CodeLengths}$: 8-bit unsigned integer arrays of size 256
  • $\mathsf{Gaps}$: 5-bit unsigned integer array (one entry per thread in each block)
  • $\mathsf{BlockOutputPos}$: 32-bit unsigned integer array (one entry per block)
  • $\mathsf{Outputs}$: BFloat16 array, for storing results
  • $B,T,n,k$: the number of thread blocks, number of threads, number of bytes processed by each thread, number of compact LUTs, respectively
3:  Divide $\mathsf{EncodedExponent}$ into chunks:
4:    $\mathsf{EncodedExponent}_{1},\dots,\mathsf{EncodedExponent}_{B}$ of size nTnT bytes each
@@
6:    Load $\mathsf{EncodedExponent}_{b}$ into SRAM
7:    Divide $\mathsf{EncodedExponent}_{b}$ into chunks:
8:      $\mathsf{EncodedExponent}_{b,1},\dots,\mathsf{EncodedExponent}_{b,T}$ of size $n$ bytes each
9:    Load $\mathsf{LUT}_{1},\dots,\mathsf{LUT}_{k},\mathsf{CodeLengths}$ into SRAM
10:    Initialize integer arrays $\mathsf{NumElements}[1\dots T],\mathsf{ThreadOutputPos}[1\dots T]$ with all 0s
11:    Initialize BFloat16 write buffer $\mathsf{WriteBuffer}$ in SRAM
12:    for all $t\leftarrow 1,\dots,T$ (in parallel across threads) do
[…]
14:     $\mathrm{BitOffset}\leftarrow\mathsf{Gaps}[bT+t]$
15:     while $\mathrm{BitOffset}<8n$ do
16:      Read the next 4 bytes of $\mathsf{EncodedExponent}_{b,t}$, starting from the $\mathrm{BitOffset}$-th bit, into $\mathrm{Byte}_{1\dots 4}$
@@
18:      $\mathrm{Exponent}\leftarrow\mathsf{LUT}_{1}[\mathrm{Byte}_{i}]$
19:      while $\mathrm{Exponent}\geq 240$ do
20:          $\triangleright$ $\mathrm{Exponent}\geq 240$ means that it is a pointer to the next LUT
21:         $i\leftarrow i+1$
22:         $\mathrm{Exponent}\leftarrow\mathsf{LUT}_{(257-\mathrm{Exponent})}[\mathrm{Byte}_{i}]$
23:      end while
24:      $\mathrm{BitOffset}\leftarrow\mathrm{BitOffset}+\mathsf{CodeLengths}[\mathrm{Exponent}]$
25:      $\mathsf{NumElements}[t]\leftarrow\mathsf{NumElements}[t]+1$
26:     end while
27:     Thread Synchronization Barrier
28:      $\triangleright$ Compute prefix-sum using Blelloch’s Algorithm:
29:     $\mathsf{ThreadOutputPos}[t]\leftarrow\mathsf{BlockOutputPos}[b]+\sum_{i=1}^{t-1}\mathsf{NumElements}[i]$
[…]
31:     $\mathrm{BitOffset}\leftarrow\mathsf{Gaps}[bT+t]$
32:     while $\mathrm{BitOffset}<8n$ do
33:      Read the next 4 bytes of $\mathsf{EncodedExponent}_{b,t}$, starting from the $\mathrm{BitOffset}$-th bit, into $\mathrm{Byte}_{1\dots 4}$
34:      $i\leftarrow 1$
35:      $\mathrm{Exponent}\leftarrow\mathsf{LUT}_{1}[\mathrm{Byte}_{i}]$
36:      while $\mathrm{Exponent}\geq 240$ do
37:          $\triangleright$ $\mathrm{Exponent}\geq 240$ means that it is a pointer to the next LUT
38:         $i\leftarrow i+1$
39:         $\mathrm{Exponent}\leftarrow\mathsf{LUT}_{(257-\mathrm{Exponent})}[\mathrm{Byte}_{i}]$
40:      end while
41:      $\mathrm{Byte}\leftarrow\mathsf{PackedSignMantissa}\big[\mathsf{ThreadOutputPos}[t]\big]$
42:      $\mathrm{Sign}\leftarrow\mathrm{Byte}\;\texttt{bitwise\_and}\;\texttt{0b10000000}$
43:      $\mathrm{Mantissa}\leftarrow\mathrm{Byte}\;\texttt{bitwise\_and}\;\texttt{0b01111111}$
44:      $\mathsf{WriteBuffer}[\mathsf{ThreadOutputPos}[t]-\mathsf{BlockOutputPos}[b]]\leftarrow$